2017B schedule, programs, and run summaries

List of 2017B programs for: MegaCam, WIRCam, ESPaDOnS, SITELLE.

QrunID Instrument HST Dates # Nights Status Run Summary
Shutdown Aug 1-5 5 Done The shutdown was completed one day earlier than our most optimistic schedule.
17BQ01 ESPaDOnS Aug 6-14 9 Done The first 3 nights were affected by clouds, but we observed nonetheless. The seeing stabilized more quickly than anticipated; only the first night had Image Quality uncharacteristically above 1 arcsec. On the second night, we were well below 1 arcsec as usual.
17BQ02 MegaCam Aug 15-30 16 Done Despite 28 hours lost to bad weather (clouds or fog), this run was very successful. Besides 20 minutes lost to a filter jukebox issue, and an hour septn investigating network and computer issues, there were no other time lost to technical issues.
17BQ03 WIRCam Aug 31 - Sep 4 5 Done One of the two WIRCam filter wheels failed on Sep 3; althought we were able to finish the night using only one filter, the issue requires warming up and opening the cryostat. The night of Sep 4/5 was lost to high humidity. ESPaDOnS will replace WIRCam for the remaining 6 nights.
17BQ04 ESPaDOnS Sep 5 - 10 6 Done This ESPaDOnS run was very successful. Although thin cirrus affected some nights, we never stopped observing.
17BQ05 MegaCam Sep 11-24 14 Done This was an excellent run for MegaCam, with only 16 hours lost to weather (mostly high humidity), and less than 1h lost to technical issues.
17BQ06 SITELLE Sep 25-28 4 Done This short SITELLE run was successful. Only 1h was lost to a technical issue with the windscreen. Except for a couple of hours with thin cirrus, the sky was photometric. The Image Quality, however, was very poor on the last night (1.5 to 2 arcsec).
17BQ07 WIRCam Sep 29 - Oct 9 11 Done The defective filter wheel parts that failed on Sep 5 were replaced and tested in time to mount WIRCam Sep 29. WIRCam replaced the ESPaDOnS run that was scheduled for Sep/Oct. Most nights were affected by either thin cirrus or thicker clouds; for 3 of 11 the nights, most of the time was lost to high extinction. With about half of the nights remaining for this semester, about half of the hours for A & B programs have been validated.
17BQ08 SITELLE Oct 10-19 10 Done This was a challeging run, with clouds or high humidity affecting a few hours every night. The first night was also lost to a power outage. On the bright side, it is possible to reduce data taken over 2 or more nights, and piece together the observations.
17BQ09 MegaCam Oct 20-26 7 Done This short MegaCam run was affected by bad weather, with 3.5 nights lost to high humidity, and a few hours of thin cirrus on other nights.
17BQ10 ESPaDOnS Oct 27 - Nov 8 13 Done A very good ESPaDOnS run where we were able to focus on high priority programs and also get observations for quite a few overfill programs.
17BQ11 SITELLE Nov 9-13 5 Done Despite the unfavorable forecast, this last SITELLE run of the semester went relatively well. One night was lost to weather, and clouds affected all other nights to various degrees.
17BQ12 MegaCam Nov 14-27 14 Done This MegaCam run was challenged by high winds (almost 3 nights lost) and adverse weather conditions, like clouds, fog, and snow (almost 4 nights lost). The other 7 nights generally saw photometric skies with good Image Quality.
17BQ13 WIRCam Nov 28 - Dec 7 10 Done This WIRCam run, especially the first 4 nights, was affected by bad weather or bad Image Quality. Only the very last night of the run had perfect conditions for observing.
17BQ14 MegaCam Dec 8-25 17 Done The weather severly impacted the productivity of this winter run. MegaCam cooling also started to fail toward the end of the run. The exchange to ESPaDOnS was performed one day early.
17BQ15 ESPaDOnS Dec 26 - Jan 9 16 Done This last run of semester 2017B was very successful. The weather started with variable clouds, precipitation, or high winds, but improved quickly. Only 15 hours were lost to weather. Only 2.5h hours were lost when a cable fell off and a network switch became unplugged. Half a night was used for engineering. Nine of the 14 A or B programs have obtained all their observations. The remaining programs have obtained between 80% and 95% of the requested observations.
17BQ16 MegaCam Jan 10-24 15 Done This last MegaCam run of the semester was challenged by bad weather (60h lost) and cooling issues with the instrument (a total of 18h lost on 3 separate nights).
17BQ17 WIRCam Jan 25-28 4 Done 2017B programs that would have been observed during the last 3 nights of January will be observed during the February WIRCam run. This short run was completely lost to bad weather (snow, ice, high humidity).
17BQ18 ESPaDOnS Jan 29-31 3 Started 2018A ESPaDOnS programs will be started during this end-of-January run, and into February.